two angles are supplementary angles if the sum of their measures is 180. Find two supplementary angles such that the measure of the first angle is 54 less than three times the measure of the second
Answer by rfer(12678)

x+3x-54=180
4x=180+54
4x=234
x=234/4
x=58.5
3x-54=121.5